OpenStudy (anonymous):

Let f(x) = 12 over the quantity of 4 x + 2. Find f(-1).

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

good thing mathematicians invented symbols is this \[f(x)=\frac{12}{4x+2}\]?

OpenStudy (anonymous):

yes it is haha

OpenStudy (misty1212):

you want \(f(-1)\)?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

where you see an \(x\) put a \(-1\) and compute \[f(-1)=\frac{12}{4\times (-1)+2}\]

OpenStudy (anonymous):

so f(-1) 12 over -4+2